OpenStudy (anonymous):

The atomic mass of iron is 55.85, and oxygen is 16.0. What is the mass of 6.02 × 1023 molecules of Fe2O3? A. 71.85 g B. 143.7 g C. 159.7 g D. 199.55 g E. 215.55 g

OpenStudy (anonymous):

6.02x10^23=1 mol molecular mass of Fe2O3: 55.85x2+16.0x3 = 159.7 Mass of 1 mol of Fe2O3=159.7x1=159.7 (g)